OpenStudy (anonymous):

How do I describe how to find the multiplier for a rate decay of 15%?

OpenStudy (mathstudent55):

(1 - r) = 1 - 15% = 1 - 0.15 = 0.85

OpenStudy (anonymous):

\[dN=Kdt\]